OpenStudy (anonymous):

I need help How do you do this? Complete the trinomial so it is a perfect square x^2+__+64

OpenStudy (anonymous):

\[\pm16x\]

OpenStudy (anonymous):

Thanks! Can you show me how to do that step by step?

OpenStudy (anonymous):

x2-16x+64= (x-8)^2 x2+16x+64=(x+8)^2

OpenStudy (anonymous):

64=8^2 so, it will be perfect square only when+/-2*8x=16x is added

OpenStudy (anonymous):

distributing medals? alright

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

u need medals